# Pseudoentanglement from tensor networks

### v2

@@ Abstract
S2 Computational indistinguishability between $\mathcal{E}_{\rm rMPS^{\prime}}$ and $\mathcal{E}_{\rm rMPS}$
Here we give a proof of Lemma 1, stating the indistinguishability between the ensemble of MPS made out of Haar-random gates, $\mathcal{E}_{\rm rMPS}$, and the analogous ensemble where the MPS is made out of random ‘PFC’ gates, $\mathcal{E}_{\rm rMPS^{\prime}}$. To this end, we use a key property of the PFC ensemble:
@@
The moment operators for the two ensembles $\mathcal{E}_{\rm rMPS^{\prime}}$, $\mathcal{E}_{\rm rMPS}$ can be written as
@@
$\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}$ $=\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}\circ\Phi^{(m)}_{\rm rPFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm rPFC}|_{[1,\nu+1]}(\rho_{0}),$ (S7)
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}-\rho^{(m)}_{\mathcal{E}_{\rm rMPS^{\prime}}}\|_{\rm tr}\leq L\epsilon,$ (S8)
@@
Here we modify the proof presented in the main text for MPS to general isometric tensor network architectures. We define the state ensembles $\mathcal{E}_{\rm pTNS}$ (made of pseudorandom PFC gates), $\mathcal{E}_{\rm rTNS^{\prime}}$ (made of random PFC gates) and $\mathcal{E}_{\rm rTNS}$ (made of Haar-random gates) in the same ways as $\mathcal{E}_{\rm pMPS}$, $\mathcal{E}_{\rm rMPS^{\prime}}$ and $\mathcal{E}_{\rm rMPS}$, just replacing the MPS architecture with a general tensor network state (TNS).
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rTNS^{\prime}}}-\rho^{(m)}_{\mathcal{E}_{\rm rTNS}}\|_{\rm tr}\leq O(n_{U}m/\sqrt{d_{\rm min}}),\qquad d_{\rm min}=\min_{u}\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}.$ (S12)
@@
Here we generalize Lemma 2 to the more general class of tensor network states (TNS) beyond MPSs. We aim to prove that the ensemble of TNSs built out of Haar-random gates is indistinguishable from the Haar-random ensemble on the whole Hilbert space. In analogy with the MPS case, we can compute the $m$-copy distinguishability, perform the Haar averages, and obtain a modified graph corresponding to the partition functions in Fig. 2(b,d). The new graph is undirected, weighted and labeled—each edge has one of two labels, that we denote by ‘solid’ and ‘dashed’ in keeping with Fig. 1 and Fig. 2. The new graph (interaction graph of the partition function) is built from the old one (the tensor network diagram) according to these rules:
@@
The proof of computational indistinguishability proceeds in the same way as for the MPS case until the derivation of Eqs. (3-4).
@@
To prove that $\Delta_{m}^{\rm u.}$ is small, Eq. (5), we need to show that the Boltzmann weight $e^{-E[\{e,e\}]}$ of a uniform spin configuration is close to $D^{-m}$. We have
@@
The derivation of Eq. (7) makes no reference to the MPS structure and thus proceeds unchanged. Once we have a partition function with positive Boltzmann weights
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm pTNS}}-\rho^{(m)}_{\mathcal{E}_{\rm Haar}}\|_{\rm tr}\leq O(n_{U}m^{2}/\chi_{\rm min})+O(n_{U}m/\sqrt{d_{\rm min}}).$ (S17)
@@
Here we prove a lower bound on the average entropy $S(A)$ of a subsystem $A$ in the pseudorandom holographic tensor network ensemble, $\mathcal{E}_{\rm holo}$, considered in the main text. Our strategy for lower-bounding $S(A)$ is to evaluate the average purity, $\mathbb{E}[{\rm Tr}(\rho_{A}^{2})]$. Indeed we have the following:

### v3

@@ Acknowledgements.
S2 Computational indistinguishability between $\mathcal{E}_{\rm rPFCMPS}$ and $\mathcal{E}_{\rm rMPS}$
Here we give a proof of Lemma 1, stating the indistinguishability between the ensemble of MPS made out of Haar-random gates, $\mathcal{E}_{\rm rMPS}$, and the analogous ensemble where the MPS is made out of random ‘PFC’ gates, $\mathcal{E}_{\rm rPFCMPS}$. To this end, we use a key property of the PFC ensemble:
@@
The moment operators for the two ensembles $\mathcal{E}_{\rm rPFCMPS}$, $\mathcal{E}_{\rm rMPS}$ can be written as
@@
$\rho^{(m)}_{\mathcal{E}_{\rm rPFCMPS}}$ $=\Phi^{(m)}_{\rm rPFC}|_{[L,L+\nu]}\circ\Phi^{(m)}_{\rm rPFC}|_{[L-1,L-1+\nu]}\circ\cdots\circ\Phi^{(m)}_{\rm rPFC}|_{[1,\nu+1]}(\rho_{0}),$ (S7)
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rMPS}}-\rho^{(m)}_{\mathcal{E}_{\rm rPFCMPS}}\|_{\rm tr}\leq L\epsilon,$ (S8)
@@
Here we modify the proof presented in the main text for MPS to general isometric tensor network architectures. We define the state ensembles $\mathcal{E}_{\rm pPFCTNS}$ (made of pseudorandom PFC gates), $\mathcal{E}_{\rm rPFCTNS}$ (made of random PFC gates) and $\mathcal{E}_{\rm rTNS}$ (made of Haar-random gates) in the same ways as $\mathcal{E}_{\rm pPFCMPS}$, $\mathcal{E}_{\rm rPFCMPS}$ and $\mathcal{E}_{\rm rMPS}$, just replacing the MPS architecture with a general tensor network state (TNS).
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rPFCTNS}}-\rho^{(m)}_{\mathcal{E}_{\rm rTNS}}\|_{\rm tr}\leq O(n_{U}m/\sqrt{d_{\rm min}}),\qquad d_{\rm min}=\min_{u}\prod_{\ell\in{\rm outgoing}(u)}\chi_{\ell}.$ (S12)
@@
Here we generalize Lemma 2 to the more general class of tensor network states (TNS) beyond MPSs. We aim to prove that the ensemble of TNSs built out of Haar-random gates is indistinguishable from the Haar-random ensemble on the whole Hilbert space. In analogy with the MPS case, we can compute the $m$-copy distinguishability, perform the Haar averages, and obtain a modified graph corresponding to the partition functions in Fig. 2(b,d) in the main text. The new graph is undirected, weighted and labeled—each edge has one of two labels, that we denote by ‘solid’ and ‘dashed’ in keeping with Fig. 1 and 2 in the main text. The new graph (interaction graph of the partition function) is built from the old one (the tensor network diagram) according to these rules:
@@
The proof of computational indistinguishability proceeds in the same way as for the MPS case until the derivation of Eqs. (3-4) in the main text.
@@
To prove that $\Delta_{m}^{\rm u.}$ is small, Eq. (5) of the main text, we need to show that the Boltzmann weight $e^{-E[\{e,e\}]}$ of a uniform spin configuration is close to $D^{-m}$. We have
@@
The derivation of Eq. (7) in the main text makes no reference to the MPS structure and thus proceeds unchanged. Once we have a partition function with positive Boltzmann weights
@@
$\|\rho^{(m)}_{\mathcal{E}_{\rm rTNS}}-\rho^{(m)}_{\mathcal{E}_{\rm Haar}}\|_{\rm tr}\leq O(n_{U}m^{2}/\chi_{\rm min})+O(n_{U}m/\sqrt{d_{\rm min}}).$ (S17)
@@
Here we prove a lower bound on the average entropy $S(A)$ of a subsystem $A$ in the pseudorandom holographic tensor network ensemble, $\mathcal{E}_{\rm holoTNS}$, considered in the main text. Our strategy for lower-bounding $S(A)$ is to evaluate the average purity, $\mathbb{E}[{\rm Tr}(\rho_{A}^{2})]$. Indeed we have the following: